OpenStudy (kamille):

There are two functions g(x) and f(x), which intersect at point A. I need to find an oordinate of that point. Functions: \[f(x)=5-\log _{2}(x-4)\] \[g(x)=\log _{2}x\] How I do: \[5-\log _{2}(x-4)=\log _{2}x\] \[-\log _{2}(x-4)-\log _{2}x=-5|:(-1)\] \[\log _{2}(\frac{ x+4 }{ x })=5\] \[\log _{2}(\frac{ x+4 }{ x })=\log _{2}32\] So I make a system of equation: \[\frac{ x+4 }{ x }=32\] \[\frac{ x+4 }{ x }>0\] \[x \neq 0\] I solve it and get that x=28. When put x=28 into functions, but get wrong answer. can anyone help me?: ) The answer should be 2.

OpenStudy (kamille):

@satellite73

OpenStudy (kamille):

@amistre64

OpenStudy (kamille):

@phi

OpenStudy (kamille):

@anybody?:o

OpenStudy (amistre64):

gonna have to take a break halfway thru .. thats alot of reading

OpenStudy (amistre64):

f(x)=5-log2(x-4) g(x)=log2x is that a log base 2?

OpenStudy (kamille):

oh, maybe you could do it by yourself and tell me if you get 2?

OpenStudy (kamille):

yes, the base is 2. I used Latex, doesnt it work?

OpenStudy (amistre64):

oh, you did dint you :) its been a long day

OpenStudy (kamille):



OpenStudy (amistre64):

\[f(x)=5-log_2(x-4)\\ g(x)=log_2(x)\] equate them to find out where they are the same, so a system of equations is fine \[5-log_2(x-4)=log_2(x)\] \[5=log_2(x)+log_2(x-4)\] \[5=log_2(x(x-4))\] \[2^5=2^{log_2(x(x-4))}\] \[2^5=x(x-4)\] \[32=x^2-4x\] \[36=x^2-4x+4\] \[36=(x-2)^2\] \[\pm6=x-2\] \[2\pm6=x=8,-4\] isolate for any ill gotten values

OpenStudy (kamille):

oh, right. it was +,so I needed to multiply. Thanks a lot!

OpenStudy (amistre64):

id go with x=8 meself .... since log(-4) is not real .....

OpenStudy (kamille):

yeah, but I need to find y(oordinate), so I will get 2, because f(2)=5-log(base 2)8=2

OpenStudy (kamille):

thank you alot:)

OpenStudy (amistre64):

log2(8) log2(2^3) 3 log2(2) = 3

OpenStudy (kamille):

well, there is 5 in the function too;) so 5-3=2 : ) You should get some rest! THANKS!

OpenStudy (amistre64):

x = 8, so find f(8) and g(8)

OpenStudy (amistre64):

lol, fine fine fine ;) good luck

OpenStudy (kamille):

good luck you too!

OpenStudy (amistre64):

log2(8-4) = log(2^2) = 2 5 - 2 = 3

OpenStudy (kamille):

oh haha, my mistake!

OpenStudy (kamille):

maybe the given answer is wrong